OpenStudy (raffle_snaffle):

A particular telephone number is used to receive both voice calls and fax messages. Suppose that 25% of the incoming calls involve fax messages, and consider a sample of 25 incoming calls. (Round your answers to three decimal places.) (a) What is the probability that at most 4 of the calls involve a fax message? (b) What is the probability that exactly 4 of the calls involve a fax message? (c) What is the probability that at least 4 of the calls involve a fax message? (d) What is the probability that more than 4 of the calls involve a fax message?

OpenStudy (raffle_snaffle):

@mathmale

OpenStudy (raffle_snaffle):

@zepdrix

OpenStudy (raffle_snaffle):

Binomial distribution?

zepdrix (zepdrix):

Oo boy I actually don't know >.< Sorry. Not familiar with probability stuff. When I first looked at it, I thought I knew, but nope :3 lol

OpenStudy (raffle_snaffle):

It's okay thanks for trying

OpenStudy (raffle_snaffle):

I found a solution online. The math is easy I am just trying to understand the concept.

zepdrix (zepdrix):

I'm reading about Binomial Distribution on wikipedia... maybe we can figure this out... https://en.wikipedia.org/wiki/Binomial_distribution We want cumulative distribution for the first part, ya? Because we have to consider 0, 1, 2, 3 or 4 fax messages coming in the 25. \[\large\rm \Pr(X\le4)=\sum_{i=0}^{4}\left(\begin{matrix}25 \\ \rm i\end{matrix}\right)p^i(1-p)^{25-i}\]Where our probability p=1/4. Although, if you've already found a solution, then maybe this just makes it more confusing XD lol

OpenStudy (raffle_snaffle):

well I am trying to find that formula on my lecture slides. there are several formulas that are very similar

OpenStudy (raffle_snaffle):

I have this one p(y) = [(n!)/y!(n-y)]*p^y*(q)^(n-y)

zepdrix (zepdrix):

Ok good, yes that's what I have in my summation, but with different letters. Small typo though, the (n-y) should be factorial.

zepdrix (zepdrix):

\[\large\rm \Pr(y)=\frac{n!}{y!(n-y)!}p^y q^{n-y}\]Where q=1-p.

OpenStudy (raffle_snaffle):

I see what q is but what is it? What does it illustrate?

OpenStudy (raffle_snaffle):

Can you just explain to me what n,y,and q represent?

zepdrix (zepdrix):

p represents the given probability of an even happening. For our problem, that's the 25%. q is the compliment of that, odds of event NOT happening. so q=75%

zepdrix (zepdrix):

n = total trials

zepdrix (zepdrix):

y is a `specific number of successful trials`. So when we write Pr(y), we're asking, what are the odds of getting exactly `y` fax messages from the 25 total calls we receive.

zepdrix (zepdrix):

So then for part B, Pr(4) gives us our answer. That make sense? That is the probability of getting `exactly` 4 faxes out of 25 calls.

OpenStudy (raffle_snaffle):

Wait.. so our sample total sample is 100 but we are only drawing 25 from the initial sample, so the compliment is 75 found by q = 100-25 = 75?

zepdrix (zepdrix):

maybe I should be more clear: p = odds of success q = odds of failure If odds of success = 25%, then odds of failure = 100% - 25% = 75% It's always based off of 100%. Sometimes you fail, sometimes you succeed, but `always`, 100% of the time, something happens, right? So success + failure = 100%

zepdrix (zepdrix):

The 25 sample had nothing to do with our p or q. It was the 25% that was our magic number.

OpenStudy (raffle_snaffle):

Okay that makes sense.

zepdrix (zepdrix):

Let's plug in what we know so far,\[\large\rm \Pr(y)=\frac{25!}{y!(25-y)!}\left(\frac{1}{4}\right)^y \left(\frac{3}{4}\right)^{25-y}\]

zepdrix (zepdrix):

I plugged in `25 for every n`, `1/4 which is equivalent to 25% for our p`, `3/4 which is equivalent to 75% for our q`.

OpenStudy (raffle_snaffle):

I get that. So we are going to sum 4 times?

zepdrix (zepdrix):

For part A, we want to know, Pr(0) `odds of getting zero faxes in total` Pr(1) `odds of getting one fax` Pr(2) `odds of getting two faxes` ... Pr(3) Pr(4) Yes, we have to add all of those up.

OpenStudy (raffle_snaffle):

Okay let me work on this for a bit. I will attach an image with my work in a bit..

zepdrix (zepdrix):

cool

OpenStudy (raffle_snaffle):

I am going to be awhile so if you want to help other people then go right ahead.

zepdrix (zepdrix):

I'm doing it on paper as well XD lol

zepdrix (zepdrix):

You have some options for shortcutting the first fraction if you like.

zepdrix (zepdrix):

\[\large\rm \frac{n!}{y!(n-y)!}=\left(\begin{matrix}n \\ \rm y\end{matrix}\right)=nCy\]

zepdrix (zepdrix):

We would read this as "n choose y". You can use your calculator if that's easier for you. For example, at least with my calculator, I can type 25 nCr 2 and it gives me 300 That's a shortcut to writing out 25!/2!(25-2)! which will be a little more burdensome.

zepdrix (zepdrix):

Whatever works for you though.

OpenStudy (raffle_snaffle):

is that a quicker way of writing it? The solution I found on line uses the formula you stated up above. I was hesitant to use it because I didn't know where it came from.

zepdrix (zepdrix):

Yes, if your teacher hasn't taught you that method yet, maybe he/she wants you to do it out the long way. Seems tedious though :P I'm using a simple TI-30XIIS. I think I have all of the graphing calculators on my phone/tablet though. I can probably find the button.

zepdrix (zepdrix):

Mmm no, I guess I don't have that fancy beast :) TI-82, 83, 85 and 86. Maybe same buttons though, if you need to know where it is.

OpenStudy (raffle_snaffle):

There has be some quicker way to solving these... I am just going to use nCy

OpenStudy (raffle_snaffle):

what is C?

OpenStudy (raffle_snaffle):

How do you use nCy?

zepdrix (zepdrix):

Oh oh, ok I think I found the buttons. You go, `2nd` `Multiply(math in yellow)` `F2 to get PROB` `F3 to get nCr`

zepdrix (zepdrix):

But before you press F3 to get your nCr, fire type `25`, then press `F3` then type `0` That will calculator (25 0) for you, the first one.

OpenStudy (raffle_snaffle):

hold on here

zepdrix (zepdrix):

It will leave the probability window up for you, so you don't have to navigate through "math" every time

zepdrix (zepdrix):

k

OpenStudy (raffle_snaffle):

25,nCr(0) ?

OpenStudy (raffle_snaffle):

I found nCr()

zepdrix (zepdrix):

Did it give you brackets? Maybe your calculator wants you to put it in like this then, nCr(25,0)

zepdrix (zepdrix):

mine is a little different maybe :P

OpenStudy (raffle_snaffle):

when I did nCr(25,1) = 1

zepdrix (zepdrix):

You mean nCr(25,0)?

OpenStudy (raffle_snaffle):

YEAH LOL

zepdrix (zepdrix):

nCr(25,1) should give you 25 D:

zepdrix (zepdrix):

ok good :)

zepdrix (zepdrix):

So we've got a handle on the shortcut, yay

OpenStudy (raffle_snaffle):

the 1 is illustrating we are starting at 1 then going to 4?

OpenStudy (raffle_snaffle):

nCr(25,1 <------ ?

zepdrix (zepdrix):

Yes, Your Pr(0) has a nCr(25,0) in it. Your Pr(1) has a nCr(25,1) in it. ... and so on. If that's what you were asking.

OpenStudy (raffle_snaffle):

yes yes that is what I am asking. Awesome

zepdrix (zepdrix):

\[\large\rm \Pr(y)=\color{orangered}{\frac{25!}{y!(25-y)!}}\left(\frac{1}{4}\right)^y \left(\frac{3}{4}\right)^{25-y}\] So this is what we've done to shorten things up for ourselves,\[\large\rm \Pr(y)=\color{orangered}{nCr(25,y)}\left(\frac{1}{4}\right)^y \left(\frac{3}{4}\right)^{25-y}\]

OpenStudy (raffle_snaffle):

thank you so much and thank you for explaining it to me. I should be fine here on out.

zepdrix (zepdrix):

Cool :) Was good for me to brush up on some of this stuff too haha

OpenStudy (raffle_snaffle):

I've got one more question

OpenStudy (raffle_snaffle):

@zepdrix

zepdrix (zepdrix):

sup :)

OpenStudy (raffle_snaffle):

instead of multiplying every nCr(25,1) by each increment can I just input this in my calculator nCr(25,4)*[P^y(1-p)^(n-y)+....]

OpenStudy (raffle_snaffle):

is that how it works?

OpenStudy (raffle_snaffle):

is there a function? How does it know r=y?

zepdrix (zepdrix):

`n` and `r` are the variables we normally use. we're using y instead of r, no big deal though.

OpenStudy (raffle_snaffle):

so everywhere there is a y i will have to use an r...

zepdrix (zepdrix):

Hmm I'm not sure what you're asking. You're asking if you can put it into your calculator as a function? Or you're asking if you can factor something out of each term?

OpenStudy (raffle_snaffle):

can i factor nCr out of each term?

zepdrix (zepdrix):

No, because they're all different :\ It's nCr(25,0) in the first term, nCr(25,1) in the next, and so on. None of those are the same.

zepdrix (zepdrix):

If this is way tedious, you can ... "super shortcut" and use wolfram :) lol

OpenStudy (raffle_snaffle):

I can't use my laptop on the exam XD

zepdrix (zepdrix):

mmm true true -_-

OpenStudy (raffle_snaffle):

I thought nCr was a function

OpenStudy (raffle_snaffle):

I will figure it out, thanks though!